OpenStudy (anonymous):

Betty is thinking of 2 consecutive integers whose sum is 41. Let x represent the smaller unknown integer. What are Betty's integers?

OpenStudy (amistre64):

id start with 41/2

OpenStudy (anonymous):

20 + 21 = 41 x + x+1 = 41 2x + 1 = 41

OpenStudy (anonymous):

x = 20